1. AJ Imperial is a self-employed professional who offers accounting and business management services. Her total gross receipts amounted to 3,500,000.00 for taxable year 2020. Her recorded cost of service and operating expenses were 1,200,000.00 and 800,000.00 respectively. How will you compute for AJ's tax liability? How much is her gross income, taxable income and graduated income tax due for the year?


1
Expert's answer
2021-12-13T10:26:49-0500

Solution:

AJ’s tax liability will be determined by calculating his gross income, which is the amount taxable.

Gross income = Total revenues – total expenses

Total revenue = 3,500,000

Total expenses = Cost of service + operating expenses = 1,200,000 + 800,000 = 3,000,000

Gross income = 3,500,000 – 3,000,000 = 500,000

Taxable income = 500,000